OpenStudy (anonymous):

What is the area of a circle with a diameter of 10 ft? Choose one answer. a. 25π ft^2 b. 20π ft^2 c. 10π ft^2 d. 100π ft^2

OpenStudy (anonymous):

radius will be diameter divided by 2 so r=5 ft A=πr^2 A=5^2π A=25π ft^2 so A is the answer